If z1 + z2 +z3 are 3 complex number in AP then they lie on
  1. circle
  2. ellipse
  3. straight line
  4. parabola

Grade:Select Grade

2 Answers

noogler
8 years ago
2Z2=Z1+Z3 
2(x2+iy2)=(x1+x3)+i(y1+y3)
comparing real and imaginary parts
x2=x1+x3/2        ,            y2=y1+y3/2
Z2 is midpoint of Z1andZ3
they lie on a st.line
 
APPROVE IF USEFUL!!!!
sadhan
6 years ago
A line makes an angle with positive directions90, 90, 90 degrees with X, Y, Z axis, then the direction cosines are ---------